OpenStudy (anonymous):

Qura currently pays $800 each month to rent her apartmen. Due to inflation, however, her rent is increasing by $50 each year. Meanwhile, her monthly take-home pay is $1500 and she predicts that her monthly pay will only increase by $15 eac hyear. Assuming htat her rent and take-home pay will continey to grow linearly, will herrent ever equal her take-home pay? If so, when? And how much will rent be that year?

OpenStudy (anonymous):

f(x) = $800 + 50x (x in years) g(x) = $1500 + 15x (x in years) f(x) = g(x) if: $800 + 50x = $1500 + 15x 35x = 700 x = 20 years f(x) = $800 + 20*50 = $1800